OpenStudy (anonymous):

There are 10 cards, labeled from 1 to 10, lying face down on a table. You pick a card, and if that card is a 10, you win $20. If the card is not 10 you win $0. What is the expected amount you would win (in dollars)?

OpenStudy (anonymous):

idk i think its 20

OpenStudy (anonymous):

bt thats to obvious right

OpenStudy (anonymous):

expected value is the amount you win times the probability you win it (with the sophistication that you can win a negative amount of money)

OpenStudy (anonymous):

in this case it is \[20\times \frac{1}{10}=2\]

OpenStudy (anonymous):

thanks I ll try it

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

you can also think like this pick all 10 cards, you win $20 and lose nothing averaged over the total number of picks you won \(\frac{20}{10}=2\) per pick

OpenStudy (anonymous):

I check I know its right thanks alot man

OpenStudy (anonymous):

yw